OpenStudy (anonymous):

Can anyone please answer this.. A box contains six number cards. The six numbers form a pattern when ordered from least to greatest. Jay draws these four cards from the box. 1/2 , 6 1/2, 1 1/4 and 1/4. What number cards are likely to be left in the box?

OpenStudy (anonymous):

I would put the cards I have in order smallest to largest seems a good start 1/4 - 1/2 - 1 1/4 - 6 1/2.....hmmmm nothing springs out 1/4 - 1/2 - 5/4 - 13/2 ....still nothing obvious .25 - .5 - 1.25 - 6.5

OpenStudy (anonymous):

the other 2 cards can be anywhere in the order.. to make a pattern... but can't think of any

OpenStudy (anonymous):

Ya i am having trouble finding a pattern too...I did find a cool website that has a database of the things: http://oeis.org but that came up blank too

OpenStudy (anonymous):

If you had a 3 and a 32.5 the pattern is first card .25 * 5 = card1.25 card .5 * 5 = 3 card 6.5 * 5 = 32.5 .25/.5/1.25/3/6.5/32.5